# Explainable Framework for Swarm Intelligence Based on Fitness Landscape Features and Machine Learning

### v2

@@ 2 ELA-based visualization for benchmarking
Selecting an effective optimizer requires prior knowledge of a problem’s landscape, as algorithmic performance is heavily influenced by its underlying structure. Traditional high-level characterizations using categorical descriptions of properties like modality, separability, and variable scaling [39] provide a starting point but exhibit critical limitations. They lack quantitative precision, require specialized expertise to interpret, often miss nuanced characteristics, and assume a comprehensive problem understanding that is rarely available in practice.
@@
This gap in understanding is exacerbated by the standard tools of the field. While benchmark suites like the BBOB versions in the COCO platform [19], the CEC function suites [58], and the diverse sets in Nevergrad [47] provide essential platforms for fair and reproducible comparison, they often perpetuate a “black-box” evaluation paradigm. A primary issue is the limited generalizability of results; algorithms that excel on these specific test functions may not perform well on real-world problems. This is frequently due to a lack of benchmark diversity, the use of non-representative baselines, and the fundamental unknown nature of the internal landscape properties that dictate problem difficulty. These limitations can lead to biased comparisons and hinder progress, as the reasons for an algorithm’s success or failure remain opaque.
@@
To overcome the limitations of both high-level classifications and black-box benchmarking, we turn to quantitative, data-driven methods like low-level features [5]. Exploratory Landscape Analysis (ELA) addresses these issues by systematically characterizing problems through sampled data. ELA extracts quantitative meta-features such as convexity, distribution of objective values, local search structure, clustering, and information content from a set of candidate solutions generated via sampling. This provides deeper, more objective insights into the properties that influence algorithm performance.
@@
Tools like the R-package flacco (and its Python counterpart, Pflacco [46]) consolidate this process, offering access to over 300 numerical features across 17 feature sets from a single initial design, a dataset of sampled points and their objective values. The package manages computational overhead by tracking the function evaluations and time required for each feature set. The system delivers essential mathematical function definitions to feature sets that need additional evaluations and supports specialized methods such as cell mapping which needs block divisions across dimensions.
@@
Beyond feature calculation, flacco offers visualization techniques to gain deeper landscape insights. The plotCellMapping [25] function, for instance, visualizes the problem landscape by decomposing the continuous decision space into cells. This method identifies key cell types: attractor cells (black), uncertain cells influenced by multiple attractors (grey), and certain cells that form basins of attraction. Arrows point from each non-attractor cell towards its attractor, with length indicating attraction probability. The interactions and distributions of these cells form the basis for another set of features and provide a powerful visual summary of landscape structure.
@@
Building on these cell mapping foundations, more advanced constructs like barrier trees [27] can be generated. In this representation, local optima (valleys) are depicted as leaves (filled circles), while ridges connecting adjacent valleys are represented by branching nodes (non-filled diamonds), providing a hierarchical model of the fitness landscape’s structure.Visualization of the cell mapping plots and barrier trees plot are shown in Tables 1-3. Similarly, as detailed in Table 4, Information Content of Fitness Sequences (ICoFiS) measures properties such as smoothness, ruggedness, and neutrality by analyzing the entropy in fitness sequences generated from random walks with variable step sizes. This approach effectively adapts discrete information-theoretic measures for continuous optimization domains, with the metrics in Table 4 serving as a direct output of this analysis. The comprehensive nature of this approach is summarized in Tables 5-6, which details feature sets such as those computed by calculate_ela_meta, calculate_ela_distribution, calculate_nbc, calculate_dispersion, and calculate_information_content used to capture the global structure, value distribution, clustering behaviour, dispersion, and informational complexity of a selected BBOB function’s landscape. Section 6 details a classification methodology where an expanded suite of ELA features (building on Tables 5 and 6) is used to predict categorical performance of the PSO algorithm, as measured by its AOCC. The remaining set of ELA features, used for the classification task, has been made publicly available on GitHub.
@@ 3.1 Swarm Behaviour Analysis on Population Topologies
In this subsection, we are presenting some important communication topologies in PSO, The Star topology establishes a centralized communication framework where every particle in the swarm maintains a direct connection to the single best solution found by any member of the population. This architecture creates a fully-connected information network where all particles simultaneously receive updates from the global best position [41, 43]. The topology’s design ensures immediate propagation of the best-known solution throughout the entire swarm, creating strong directional guidance for all particles’ movement vectors. The structural configuration inherently prioritizes the dissemination of elite solutions while maintaining uniform influence across all swarm members. This complete connectivity pattern represents the most information-rich network configuration in swarm intelligence systems. The Ring topology organizes particles in a circular network where each individual maintains connections only to its immediate $k$ nearest neighbors. This decentralized architecture supports two operational modes: a static configuration with fixed neighborhood relationships, or a dynamic variant where connections may adapt during the search process. Particle interactions are governed by Minkowski $p$-norm distance metrics, with $p=1$ implementing Manhattan distance ($L_{1}$) and $p=2$ yielding Euclidean distance ($L_{2}$) calculations between swarm members. The constrained connectivity pattern creates localized information channels that naturally limit solution homogenization across the population [32, 52]. The Von Neumann topology organizes particles in a multidimensional grid configuration, where connectivity follows Delannoy number relationships. These combinatorial patterns determine each particle’s neighborhood size as a function of both the specified interaction range ($r$) and the problem’s dimensionality. The resulting lattice structure creates regular, spatially-distributed connections between particles, establishing a systematic information flow pattern across the search space. This geometric arrangement provides distinct advantages in maintaining organized particle interactions while preserving structured exploration capabilities [56, 37].
@@ 3.2 Explainer Framework for PSO
In this section, we propose an explainer framework for PSO built upon the IOHxplainer [55] environment. IOHxplainer provides a robust foundation for capturing, visualising, and quantifying algorithm dynamics by linking performance trajectories with exploratory landscape analysis (ELA) features and other descriptive measures. By extending this framework to PSO, we aim to uncover interpretable insights into swarm behaviour, parameter sensitivities, and landscape-driven performance variations. Such a framework will not only enhance transparency in understanding PSO but will also support principled algorithm selection, parameter tuning, and the design of adaptive strategies for complex optimization tasks.
@@ 3.2.1 IOHxplainer
IOHxplainer is a benchmarking platform designed to assess the performance of iterative optimization heuristics (IOHs), such as evolutionary algorithms and swarm-based algorithms [55]. It leverages explainable AI (XAI) to automatically generate insightful visualizations and statistical analyses, streamlining the evaluation of millions of algorithm configurations from large-scale empirical studies. This method helps researchers better understand how different algorithm parts work together and affect performance in various problem areas. The system collects performance data, including anytime performance measures, by testing many different algorithm setups on a range of benchmark functions. To use IOHxplainer, users start by defining the configuration space for the algorithm they want to study. This means setting value limits for continuous, integer, and categorical parameters and laying out any conditional relationships between them. The platform supports both thorough grid searches and random sampling, making it easier to explore and analyze potentially large configuration spaces. Although IOHxplainer is especially advantageous for examining modular algorithms with numerous configuration possibilities, it is also suitable for studying conventional algorithms along with their hyperparameter spaces. The configuration space is defined in a flexible manner, similar to approaches like SMAC  [16, 31], ensuring adaptability to different optimization scenarios.
@@
In this research, we utilized the IOHxplainer framework, originally designed and tested for evolutionary algorithms such as modCMA [28] and modDE [54]. We have re-invented this framework to analyze and elucidate the behavior of PSO algorithm, specifically to analyse the dynamics of PSO integrated with various communication topologies. A detailed representation of this framework is depicted in Figure 1.
@@
These enhancements aim to clarify the effects of topology configurations on the algorithm’s performance by employing a suite of XAI techniques. This enables the extraction of insightful visualizations and statistical data from detailed empirical analyses which facilitates the evaluation of potentially innumerable PSO algorithmic frameworks. It also allows for the definition and consideration of hyper-parameter dependencies, providing a nuanced approach to the configuration of the PSO algorithm and its considered topologies. For each PSO configuration, we establish a specific evaluation budget and dimensionality to test against the BBOB benchmark suite [20], which features 24 diverse noiseless functions as part of the COCO framework [18].
Upon completing the experimental runs, which are executed in parallel, the framework captures detailed performance metrics based on pre-set budget, pre-set target, or flexible-time performance measures. The ‘explainer’ component of the framework then applies advanced XAI methodologies, such as the TreeSHAP technique [35], to determine the SHAP values for components of the PSO algorithm and associated topology hyperparameters  [36]. These values indicate the marginal impact of each element on the performance metrics across different runs. Although typically approximated due to computational constraints, the SHAP values are meticulously aggregated and visualized for each function within the benchmark suite. This visualization, often presented as a swarm plot, provides PSO designers and practitioners with a clear view of how specific hyperparameters influence performance under varied conditions and across different topologies. This data, once gathered, is processed either in combination with previous datasets or as a stand-alone analysis to extract SHAP values. The details of experiments and results are compiled in subsequent sections and related resources are provided at GitHub link: https://github.com/GitNitin02/ioh_pso.
@@ 4.1 Performance Metrics
To evaluate the generated algorithms effectively over a complete set of benchmark functions, we use an anytime performance measure. This approach is critical because it captures the entire convergence profile of an algorithm, thereby rewarding those that find good solutions quickly and reliably, not just those that converge to a decent solution by the final budget. Specifically, it quantifies performance over the complete budget, instead of only looking at the final objective function value. We use the normalized Area Over the Convergence Curve (AOCC) (Eq. (3)), as introduced in  [55]. The AOCC is calculated as:
@@
AOCC metric is calculated using the series of best-found function values ($\bar{y}$) within a given evaluation budget ($B$), where the function value range is bounded between $lb$ (lower bound) and $ub$ (upper bound). This performance measure builds upon the empirical cumulative distribution function (ECDF), [34] which characterizes the probability distribution of optimization results. Specifically, AOCC represents the integrated area beneath the ECDF curve across an infinite set of target values within the specified bounds.
@@ 5 Hyperparameter Impact Analysis
We examined $1,728$ PSO configurations across twenty four BBOB functions, focusing on the impact of different communication topologies and hyperparameters on performance. Using SHAP-based plots for 2d & 5d (Table 7), we analyse the contribution parameters such as n_particles, $c_{1}$, $c_{2}$, $w$, $p$, and $k$. The hyperparameters for the PSO algorithm were selected based on established conventions and empirical ranges from the literature. The cognitive ($c_{1}$) and social ($c_{2}$) coefficients were chosen from intervals that adhere to the common stability criterion of ($c_{1}+c_{2}<4$), which helps prevent particle velocity divergence and promotes effective convergence, as analyzed by [10]. The inertia weight ($w$) values of {0.9, 0.5, 0.7} cover standard strategies, where a high value (0.9) encourages global exploration by maintaining momentum, and lower values (0.5, 0.7) facilitate local exploitation, following the principles introduced by [48]. Swarm sizes ($n$) of {50, 100, 150} were selected as a representative range commonly used in the literature, as surveyed by [6]. Finally, the parameters for neighborhood topology ($k,p,r$) were limited to low integer values to model fundamental local interaction structures—such as the Ring and Von Neumann topologies identified as high-performing by [24]—without inducing excessive computational complexity. The SHAP values indicate how each parameter affects the AOCC, with yellow dots representing positive contributions and violet dots indicating negative ones. There is a clear distinction between the two colours across all topologies and parameters, such as n_particles, $c_{1}$, $c_{2}$, $p$, $k$, etc. and others. Specifically, yellow dots represent positive SHAP values, while violet dots consistently indicate negative SHAP values.
@@ 5.3 PSO Variant configuration selection
In the context of the PSO topology comparison, $R^{2}$ (R-squared) [40] could serve as a supplementary metric to evaluate how well each topology’s performance aligns with an ideal optimization trajectory or expected convergence behaviour. For example, if the Star topology’s fast convergence in unimodal functions consistently follows a predictable pattern, its $R^{2}$ value would be high, indicating that its performance is tightly correlated with the problem’s simplicity. Conversely, in multi-modal functions, a lower $R^{2}$ for the Star topology might reflect its erratic performance due to premature convergence, whereas the Ring or Von Neumann topologies with their better diversity maintenance could exhibit higher $R^{2}$ values, suggesting more stable alignment with the problem’s complexity.
@@ 6 Data-Driven and Interpretable Configuration Learning
In the IOHxplainer pipeline, extensive data is gathered on the behaviour and performance of various algorithm configurations for each topologies. While explainable AI (XAI) [11] techniques can extract key insights from this data, there are further advantages to systematically storing and re-using it. In this section, we highlight several practical ways in which the collected data can be used to enhance the selection, configuration, and understanding of algorithm configurations.
@@
The IOHxplainer toolbox offers a valuable secondary benefit: it enables the straightforward training of machine learning models that can predict the most suitable algorithm configuration including both the choice of modules and their hyper-parameters based on observed characteristics of the optimization landscape. Figure 8 represents the workflow for performance classification using Exploratory Landscape Analysis (ELA) features. The diagram illustrates the process from feature extraction and augmentation to model training and evaluation for predicting PSO algorithm performance based on AOCC metrics. This is essentially the task of automated algorithm configuration and selection, echoing the goals of previous research such as those referenced in [26, 33] and [53]. What distinguishes our approach is that we merge the selection of the algorithm and the tuning of its configuration into a single, unified machine learning problem. We achieve this by employing multi-output models that can handle both classification (for discrete choices like module selection) and regression (for continuous settings like hyper-parameters) simultaneously. Specifically, our models are restricted to shallow Decision Tree (DT) [51, 57] and Random Forest (RF) [44]. This choice was strategic, balancing the competing needs of interpretability and predictive performance. The shallow DT was chosen for its role as a transparent, ‘white-box’ model; its limited depth yields a simple set of human-readable decision rules that directly link Exploratory Landscape Analysis (ELA) features to algorithm performance, providing crucial insights for domain experts. To complement this and ensure high accuracy, we employed a RF model. As an ensemble method, RF mitigates the potential overfitting of individual trees and robustly captures the complex, non-linear relationships between the wide range of ELA features and the optimal algorithm configuration. We gathered a wide range of Exploratory Landscape Analysis (ELA) features to show what we can do. To create these features, we set up an experiment where we sampled 1000 points from each instance in the BBOB suite. For each instance, we aimed for our machine learning models to find the best-performing algorithm configuration from our experimental results. Using this two-pronged modeling approach-RF for powerful predictions and DT for transparent rule extraction-we can train our systems to predict which algorithm configurations will perform well based on the ELA features. This approach not only automates the process of algorithm selection and configuration, but also provides transparent and understandable decision rules thanks to the interpretable nature of the models used.
@@
The AOCC performance of the Random Forest (RF) [44], Decision Tree (DT) [51], Single Best (SB), and Average Best (AB) configurations across all topologies and functions is presented in Figures 9 and 10. The key distinction is the validation strategy: Figure 9 tests generalization to unseen functions, and Figure 10 tests generalization to unseen problem instances. In Figures 11- 13, we observe how different PSO setup strategies work with Star, Ring, and Von Neumann patterns when applied to BBOB functions with d=5d{=}5, using both leave one function out (LoFo) and leave one instance out (LoIo) methods for validation. When looking at the Star topology, it’s clear that the RF model does much better than the others, achieving the lowest AOCC, whereas AB falls behind, showing the advantages of model-guided configurations. On the other hand, in the Ring topology, all three methods, AB, DT, and RF, perform similarly, with AB being a bit more stable and even holding its own against the model-based approaches. In the Von Neumann setup, AB consistently achieves the lowest AOCC, outperforming both DT and RF, which tend to be more variable. When it comes to validation methods, using LoIo often results in tighter and steadier AOCC distributions for every topology. This indicates that it does a better job of predicting how the model will perform with new instances. On the other hand, LoFo presents a tougher challenge for generalization and shows clearer differences in performance among the various methods. In general, model-based approaches seem to work best in Star topology, while AB tends to perform just as well or even better in more organized topologies like Ring and Von Neumann.

### v3

@@ 2 ELA-based visualization for benchmarking
Selecting an effective optimizer requires prior knowledge of a problem’s landscape, as algorithmic performance is heavily influenced by its underlying structure. Traditional high-level characterizations using categorical descriptions of properties like modality, separability, and variable scaling [mersmann2011exploratory] provide a starting point but exhibit critical limitations. They lack quantitative precision, require specialized expertise to interpret, often miss nuanced characteristics, and assume a comprehensive problem understanding that is rarely available in practice.
@@
This gap in understanding is exacerbated by the standard tools of the field. While benchmark suites like the BBOB versions in the COCO platform [hansen2021coco], the CEC function suites [wu2017problem], and the diverse sets in Nevergrad [rapin2018nevergrad] provide essential platforms for fair and reproducible comparison, they often perpetuate a “black-box” evaluation paradigm. A primary issue is the limited generalizability of results; algorithms that excel on these specific test functions may not perform well on real-world problems. This is frequently due to a lack of benchmark diversity, the use of non-representative baselines, and the fundamental unknown nature of the internal landscape properties that dictate problem difficulty. These limitations can lead to biased comparisons and hinder progress, as the reasons for an algorithm’s success or failure remain opaque.
@@
To overcome the limitations of both high-level classifications and black-box benchmarking, we turn to quantitative, data-driven methods like low-level features [bischl2012algorithm]. Exploratory Landscape Analysis (ELA) addresses these issues by systematically characterizing problems through sampled data. ELA extracts quantitative meta-features such as convexity, distribution of objective values, local search structure, clustering, and information content from a set of candidate solutions generated via sampling. This provides deeper, more objective insights into the properties that influence algorithm performance.
@@
Tools like the R-package flacco (and its Python counterpart, Pflacco [prager2024pflacco]) consolidate this process, offering access to over 300 numerical features across 17 feature sets from a single initial design, a dataset of sampled points and their objective values. The package manages computational overhead by tracking the function evaluations and time required for each feature set. The system delivers essential mathematical function definitions to feature sets that need additional evaluations and supports specialized methods such as cell mapping which needs block divisions across dimensions.
@@
Beyond feature calculation, flacco offers visualization techniques to gain deeper landscape insights. The plotCellMapping [kerschke2014cell] function, for instance, visualizes the problem landscape by decomposing the continuous decision space into cells. This method identifies key cell types: attractor cells (black), uncertain cells influenced by multiple attractors (grey), and certain cells that form basins of attraction. Arrows point from each non-attractor cell towards its attractor, with length indicating attraction probability. The interactions and distributions of these cells form the basis for another set of features and provide a powerful visual summary of landscape structure.
@@
Building on these cell mapping foundations, more advanced constructs like barrier trees [kerschke2019comprehensive] can be generated. In this representation, local optima (valleys) are depicted as leaves (filled circles), while ridges connecting adjacent valleys are represented by branching nodes (non-filled diamonds), providing a hierarchical model of the fitness landscape’s structure.Visualization of the cell mapping plots and barrier trees plot are shown in Tables 1-3. Similarly, as detailed in Table 4, Information Content of Fitness Sequences (ICoFiS) measures properties such as smoothness, ruggedness, and neutrality by analyzing the entropy in fitness sequences generated from random walks with variable step sizes. This approach effectively adapts discrete information-theoretic measures for continuous optimization domains, with the metrics in Table 4 serving as a direct output of this analysis. The comprehensive nature of this approach is summarized in Tables 5-6, which details feature sets such as those computed by calculate_ela_meta, calculate_ela_distribution, calculate_nbc, calculate_dispersion, and calculate_information_content used to capture the global structure, value distribution, clustering behaviour, dispersion, and informational complexity of a selected BBOB function’s landscape. Section 6 details a classification methodology where an expanded suite of ELA features (building on Tables 5 and 6) is used to predict categorical performance of the PSO algorithm, as measured by its AOCC. The remaining set of ELA features, used for the classification task, has been made publicly available on GitHub.
@@ 3.1 Swarm Behaviour Analysis on Population Topologies
In this subsection, we are presenting some important communication topologies in PSO, The Star topology establishes a centralized communication framework where every particle in the swarm maintains a direct connection to the single best solution found by any member of the population. This architecture creates a fully-connected information network where all particles simultaneously receive updates from the global best position [miranda2008stochastic, ni2013new]. The topology’s design ensures immediate propagation of the best-known solution throughout the entire swarm, creating strong directional guidance for all particles’ movement vectors. The structural configuration inherently prioritizes the dissemination of elite solutions while maintaining uniform influence across all swarm members. This complete connectivity pattern represents the most information-rich network configuration in swarm intelligence systems. The Ring topology organizes particles in a circular network where each individual maintains connections only to its immediate $k$ nearest neighbors. This decentralized architecture supports two operational modes: a static configuration with fixed neighborhood relationships, or a dynamic variant where connections may adapt during the search process. Particle interactions are governed by Minkowski $p$-norm distance metrics, with $p=1$ implementing Manhattan distance ($L_{1}$) and $p=2$ yielding Euclidean distance ($L_{2}$) calculations between swarm members. The constrained connectivity pattern creates localized information channels that naturally limit solution homogenization across the population [liu2016topology, sun2023particle]. The Von Neumann topology organizes particles in a multidimensional grid configuration, where connectivity follows Delannoy number relationships. These combinatorial patterns determine each particle’s neighborhood size as a function of both the specified interaction range ($r$) and the problem’s dimensionality. The resulting lattice structure creates regular, spatially-distributed connections between particles, establishing a systematic information flow pattern across the search space. This geometric arrangement provides distinct advantages in maintaining organized particle interactions while preserving structured exploration capabilities [von1935complete, lynn2018population].
@@ 3.2 Explainer Framework for PSO
In this section, we propose an explainer framework for PSO built upon the IOHxplainer [van2025explainable] environment. IOHxplainer provides a robust foundation for capturing, visualising, and quantifying algorithm dynamics by linking performance trajectories with exploratory landscape analysis (ELA) features and other descriptive measures. By extending this framework to PSO, we aim to uncover interpretable insights into swarm behaviour, parameter sensitivities, and landscape-driven performance variations. Such a framework will not only enhance transparency in understanding PSO but will also support principled algorithm selection, parameter tuning, and the design of adaptive strategies for complex optimization tasks.
@@ 3.2.1 IOHxplainer
IOHxplainer is a benchmarking platform designed to assess the performance of iterative optimization heuristics (IOHs), such as evolutionary algorithms and swarm-based algorithms [van2025explainable]. It leverages explainable AI (XAI) to automatically generate insightful visualizations and statistical analyses, streamlining the evaluation of millions of algorithm configurations from large-scale empirical studies. This method helps researchers better understand how different algorithm parts work together and affect performance in various problem areas. The system collects performance data, including anytime performance measures, by testing many different algorithm setups on a range of benchmark functions. To use IOHxplainer, users start by defining the configuration space for the algorithm they want to study. This means setting value limits for continuous, integer, and categorical parameters and laying out any conditional relationships between them. The platform supports both thorough grid searches and random sampling, making it easier to explore and analyze potentially large configuration spaces. Although IOHxplainer is especially advantageous for examining modular algorithms with numerous configuration possibilities, it is also suitable for studying conventional algorithms along with their hyperparameter spaces. The configuration space is defined in a flexible manner, similar to approaches like SMAC  [garcia2023explainable, lindauer2019boah], ensuring adaptability to different optimization scenarios.
@@
In this research, we utilized the IOHxplainer framework, originally designed and tested for evolutionary algorithms such as modCMA [kostovska2022importance] and modDE [van2024explainable]. We have re-invented this framework to analyze and elucidate the behavior of PSO algorithm, specifically to analyse the dynamics of PSO integrated with various communication topologies. A detailed representation of this framework is depicted in Figure 1.
@@
These enhancements aim to clarify the effects of topology configurations on the algorithm’s performance by employing a suite of XAI techniques. This enables the extraction of insightful visualizations and statistical data from detailed empirical analyses which facilitates the evaluation of potentially innumerable PSO algorithmic frameworks. It also allows for the definition and consideration of hyper-parameter dependencies, providing a nuanced approach to the configuration of the PSO algorithm and its considered topologies. For each PSO configuration, we establish a specific evaluation budget and dimensionality to test against the BBOB benchmark suite [hansen2009real], which features 24 diverse noiseless functions as part of the COCO framework [hansen2022anytime].
Upon completing the experimental runs, which are executed in parallel, the framework captures detailed performance metrics based on pre-set budget, pre-set target, or flexible-time performance measures. The ‘explainer’ component of the framework then applies advanced XAI methodologies, such as the TreeSHAP technique [lundberg2020local], to determine the SHAP values for components of the PSO algorithm and associated topology hyperparameters  [lundberg2017unified]. These values indicate the marginal impact of each element on the performance metrics across different runs. Although typically approximated due to computational constraints, the SHAP values are meticulously aggregated and visualized for each function within the benchmark suite. This visualization, often presented as a swarm plot, provides PSO designers and practitioners with a clear view of how specific hyperparameters influence performance under varied conditions and across different topologies. This data, once gathered, is processed either in combination with previous datasets or as a stand-alone analysis to extract SHAP values. The details of experiments and results are compiled in subsequent sections and related resources are provided at GitHub link: https://github.com/GitNitin02/ioh_pso.
@@ 4.1 Performance Metrics
To evaluate the generated algorithms effectively over a complete set of benchmark functions, we use an anytime performance measure. This approach is critical because it captures the entire convergence profile of an algorithm, thereby rewarding those that find good solutions quickly and reliably, not just those that converge to a decent solution by the final budget. Specifically, it quantifies performance over the complete budget, instead of only looking at the final objective function value. We use the normalized Area Over the Convergence Curve (AOCC) (Eq. (3)), as introduced in  [van2025explainable]. The AOCC is calculated as:
@@
AOCC metric is calculated using the series of best-found function values ($\bar{y}$) within a given evaluation budget ($B$), where the function value range is bounded between $lb$ (lower bound) and $ub$ (upper bound). This performance measure builds upon the empirical cumulative distribution function (ECDF), [lopez2024using] which characterizes the probability distribution of optimization results. Specifically, AOCC represents the integrated area beneath the ECDF curve across an infinite set of target values within the specified bounds.
@@ 5 Hyperparameter Impact Analysis
We examined $1,728$ PSO configurations across twenty four BBOB functions, focusing on the impact of different communication topologies and hyperparameters on performance. Using SHAP-based plots for 2d & 5d (Table 7), we analyse the contribution parameters such as n_particles, $c_{1}$, $c_{2}$, $w$, $p$, and $k$. The hyperparameters for the PSO algorithm were selected based on established conventions and empirical ranges from the literature. The cognitive ($c_{1}$) and social ($c_{2}$) coefficients were chosen from intervals that adhere to the common stability criterion of ($c_{1}+c_{2}<4$), which helps prevent particle velocity divergence and promotes effective convergence, as analyzed by [985692]. The inertia weight ($w$) values of {0.9, 0.5, 0.7} cover standard strategies, where a high value (0.9) encourages global exploration by maintaining momentum, and lower values (0.5, 0.7) facilitate local exploitation, following the principles introduced by [shi1998modified]. Swarm sizes ($n$) of {50, 100, 150} were selected as a representative range commonly used in the literature, as surveyed by [blackwell2007particle]. Finally, the parameters for neighborhood topology ($k,p,r$) were limited to low integer values to model fundamental local interaction structures—such as the Ring and Von Neumann topologies identified as high-performing by [kennedy2002population]—without inducing excessive computational complexity. The SHAP values indicate how each parameter affects the AOCC, with yellow dots representing positive contributions and violet dots indicating negative ones. There is a clear distinction between the two colours across all topologies and parameters, such as n_particles, $c_{1}$, $c_{2}$, $p$, $k$, etc. and others. Specifically, yellow dots represent positive SHAP values, while violet dots consistently indicate negative SHAP values.
@@ 5.3 PSO Variant configuration selection
In the context of the PSO topology comparison, $R^{2}$ (R-squared) [miles2005r] could serve as a supplementary metric to evaluate how well each topology’s performance aligns with an ideal optimization trajectory or expected convergence behaviour. For example, if the Star topology’s fast convergence in unimodal functions consistently follows a predictable pattern, its $R^{2}$ value would be high, indicating that its performance is tightly correlated with the problem’s simplicity. Conversely, in multi-modal functions, a lower $R^{2}$ for the Star topology might reflect its erratic performance due to premature convergence, whereas the Ring or Von Neumann topologies with their better diversity maintenance could exhibit higher $R^{2}$ values, suggesting more stable alignment with the problem’s complexity.
@@ 6 Data-Driven and Interpretable Configuration Learning
In the IOHxplainer pipeline, extensive data is gathered on the behaviour and performance of various algorithm configurations for each topologies. While explainable AI (XAI) [ding2022explainability] techniques can extract key insights from this data, there are further advantages to systematically storing and re-using it. In this section, we highlight several practical ways in which the collected data can be used to enhance the selection, configuration, and understanding of algorithm configurations.
@@
The IOHxplainer toolbox offers a valuable secondary benefit: it enables the straightforward training of machine learning models that can predict the most suitable algorithm configuration including both the choice of modules and their hyper-parameters based on observed characteristics of the optimization landscape. Figure 8 represents the workflow for performance classification using Exploratory Landscape Analysis (ELA) features. The diagram illustrates the process from feature extraction and augmentation to model training and evaluation for predicting PSO algorithm performance based on AOCC metrics. This is essentially the task of automated algorithm configuration and selection, echoing the goals of previous research such as those referenced in [kerschke2019automated, long2022learning] and [trajanov2022improving]. What distinguishes our approach is that we merge the selection of the algorithm and the tuning of its configuration into a single, unified machine learning problem. We achieve this by employing multi-output models that can handle both classification (for discrete choices like module selection) and regression (for continuous settings like hyper-parameters) simultaneously. Specifically, our models are restricted to shallow Decision Tree (DT) [song2015decision, wang2025neighborhood] and Random Forest (RF) [pal2005random]. This choice was strategic, balancing the competing needs of interpretability and predictive performance. The shallow DT was chosen for its role as a transparent, ‘white-box’ model; its limited depth yields a simple set of human-readable decision rules that directly link Exploratory Landscape Analysis (ELA) features to algorithm performance, providing crucial insights for domain experts. To complement this and ensure high accuracy, we employed a RF model. As an ensemble method, RF mitigates the potential overfitting of individual trees and robustly captures the complex, non-linear relationships between the wide range of ELA features and the optimal algorithm configuration. We gathered a wide range of Exploratory Landscape Analysis (ELA) features to show what we can do. To create these features, we set up an experiment where we sampled 1000 points from each instance in the BBOB suite. For each instance, we aimed for our machine learning models to find the best-performing algorithm configuration from our experimental results. Using this two-pronged modeling approach-RF for powerful predictions and DT for transparent rule extraction-we can train our systems to predict which algorithm configurations will perform well based on the ELA features. This approach not only automates the process of algorithm selection and configuration, but also provides transparent and understandable decision rules thanks to the interpretable nature of the models used.
@@
The AOCC performance of the Random Forest (RF) [pal2005random], Decision Tree (DT) [song2015decision], Single Best (SB), and Average Best (AB) configurations across all topologies and functions is presented in Figures 9 and 10. The key distinction is the validation strategy: Figure 9 tests generalization to unseen functions, and Figure 10 tests generalization to unseen problem instances. In Figures 11- 13, we observe how different PSO setup strategies work with Star, Ring, and Von Neumann patterns when applied to BBOB functions with d=5d{=}5, using both leave one function out (LoFo) and leave one instance out (LoIo) methods for validation. When looking at the Star topology, it’s clear that the RF model does much better than the others, achieving the lowest AOCC, whereas AB falls behind, showing the advantages of model-guided configurations. On the other hand, in the Ring topology, all three methods, AB, DT, and RF, perform similarly, with AB being a bit more stable and even holding its own against the model-based approaches. In the Von Neumann setup, AB consistently achieves the lowest AOCC, outperforming both DT and RF, which tend to be more variable. When it comes to validation methods, using LoIo often results in tighter and steadier AOCC distributions for every topology. This indicates that it does a better job of predicting how the model will perform with new instances. On the other hand, LoFo presents a tougher challenge for generalization and shows clearer differences in performance among the various methods. In general, model-based approaches seem to work best in Star topology, while AB tends to perform just as well or even better in more organized topologies like Ring and Von Neumann.